the perimeter of a rectangle is 44 cm. one side is 12 cm longer than the other side. find the lengths of the sides.
Answer by blwinbbbles(106)

l = length w = width
Let's write what we know
l = w + 12 ....one side is 12 cm longer that the other
2l + 2w = 44 ....since you add up all sides for the perimeter and there are 2 lengths and 2 widths
now use substitution since we know l = w + 12
2(w + 12) + 2w =44
2w + 24 + 2w = 44
4w + 24 = 44 subtract 24 from both sides
4w = 20 divide by 4
w = 5
now that you know what w is:
l = 5 + 12 (first equation)
l = 17 so now you know
w = 5cm and l = 17cm